OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 8 4 6 Line 2 x y –1 1 3 5 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

@pooja195

OpenStudy (aihberkhan):

Okay. In this problem we are trying to find an equation of form \(y=ax+b\), where \(a\) is slope, and \(b\) is intercept, which passes through points \((x1, y1) = (-4, 8)\) and \((x2, y2) = (4, 6)\)

OpenStudy (aihberkhan):

So, slope \(a\) is: \[a = \frac{ y2 - y1 }{ x2 - x1 } = \frac{ 6 - 8 }{ 4 - (-4) } = -0.25\]

OpenStudy (aihberkhan):

The intercept is found from this equation: \(a*x1+b = y1\) or also... \(-0.25 * (-4) + b = 7\)

OpenStudy (anonymous):

so d

OpenStudy (aihberkhan):

Nope... lets try and solve then I can help you out :)

OpenStudy (aihberkhan):

So, lets find intercept b

OpenStudy (aihberkhan):

We would find this by \(b = y1 - a * x1\) or also \(b = 8 - (-0.25) * (-4) = 7\)

OpenStudy (aihberkhan):

So the equation is \(y = (-0.25)x + 7\)

OpenStudy (aihberkhan):

Now thats for line 1. Lets do line 2 okay? :)

OpenStudy (aihberkhan):

So... now line 2... Slope \(a\) is \(1\) after solving the same way as we did for line 1. Intercept \(b\) is \(2\) solving the same way... once again. So your equation is \(y = 1x + 2\)

OpenStudy (anonymous):

so b

OpenStudy (aihberkhan):

Now solve for \(x\). You get 4 and for \(y\) you get 6. When you graph the two lines they are intersecting at point \((4, 6)\). Right! So it is B!

OpenStudy (aihberkhan):

Awesome Job! :)

OpenStudy (anonymous):

thanks

OpenStudy (aihberkhan):

No problem! @michelleball477

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @michelleball477